OpenStudy (anonymous):

If a line goes through points (-2, 9) and (0, 3), what is its slope?

OpenStudy (megannicole51):

do u know the slope formula?

OpenStudy (anonymous):

yes

OpenStudy (megannicole51):

what is it?:)

OpenStudy (anonymous):

m=rise/run...?

OpenStudy (megannicole51):

thats part of it...but if you have a set of points the formula for slope=m which equals \[\frac{ Y2-Y1 }{ X2-X1 }\]

OpenStudy (megannicole51):

out of your coordinates (-2,9) and (0,3) which is Y2 and Y1?

OpenStudy (anonymous):

i knew about that one to. I suck at algebra.

OpenStudy (anonymous):

y2 is 3 and y1 is 9...?

OpenStudy (megannicole51):

algebra, and especially slope just takes a lot of practice. it took me a long time to finally understand slope because there are a lot of different formulas to remember!

OpenStudy (megannicole51):

perfect!!!

OpenStudy (megannicole51):

and what's X2 and X1?

OpenStudy (anonymous):

One would think 4 yrs of high school would be enough.

OpenStudy (anonymous):

x2 is 0 and x1 is -2

OpenStudy (megannicole51):

correct! so now just put those into the formula and subtract and you will have your slope (m) :)

OpenStudy (anonymous):

what about simplifying?

OpenStudy (megannicole51):

well what do you end up with after you plug them in and subtract?

OpenStudy (anonymous):

simplify...?

OpenStudy (megannicole51):

you said for Y2=3 and Y1=9 so 3-9 is? and you said X2=0 and X1=-2 so 0-(-2) is?

OpenStudy (megannicole51):

we need to get to the fraction part of the answer before we simplify right!